Grade 11Physical Chemistry

A MIXTURE OF H2C2O4 AND NAHC2O4 WEIGHING 2.02G WAS DISSOLVED IN WATER AND THE SOLUTION MADE UPTO ONE LITRE .TEN MILLILITRE OF THE SOLUTION REQUIRED 3.0ML OF 0.1N SODIUM HYDROXIDE SOLUTION FOR COMPLETE NEUTRALIZATION . IN OTHER EXPERIMENT, 10.0 ML OF THE SAME SOLUTION IN HOT DILUTE SULPHURIC ACID MEDIUM, REQUIRED 4.0 ML OF 0.1N POTASSIUM PERMANGANATE SOLUTIONS FOR COMPLETE REACTION. CALCULATE THE AMOUNT OF NAH2C2O4

AYUSH
9 Years agoGrade 11

1 Answer

Sujit Kumar
8 Years ago

 

Let x be the mass of H2C2O4 in the mixture, then (2.02 x) is the mass of NaHC2O4 Amount of H2C2O4 = x/ 90gmol-1
Amount of NaHC2O4 = (2.02 x)/112gmol-1
The mixture is dissolved to make 1 liter of solution

Hence molarity of H2C2O4 = ( x/90 g) molL-1

Molarity of NaHC2O4 = ((2.02 x) / 112g ) molL-1

Reaction, H2C2O4 + 2 NaOH = Na2C2O4 +2H2O

NaHC2O4 + NaOH = Na2C2O4 +H2O

Hence, 1 mol H2C2O4 = 2eq H2C2O4

1 mol NaHC2O4 = 1eq NaHC2O4

Normality of H2C2O4 = (x/90)(2) = (x/45)

Normality of NaHC2O4 = (2.02 x/112) (1) 

Total Normality of the solution

NTotal = [(x/45) + (2.02 x/112) ]
Since 10 ml of the solution is titrated, hence the amount (in equivalent) of the mixture in 10ml of the solution will be

10 /1000 [ (x/45) + (2.02 x/112) ] = 10-2 [ (x/45) + (2.02 –x/112) ] eq......... 1

Since 3ml of 0.1N NaOH is consumed, the amount (in equivalent) of the mixture will be 3 ml / 1000 (0.1) = 3 x 10-4 eq ..........2
Equating 1 and 2 we get
3 x 10
-4 = 10-2 [ (x/45) + (2.02 x /112) ]

Solving x we get, x = 0.9g ( Oxalic acid) 2.02 x = 2.02 -0.9

= 1.12 g (NaHC2O4